# Group-Aware Partial Model Merging for Children’s Automatic Speech Recognition

### v1

@@ Abstract
Automatic Speech Recognition (ASR) for children remains challenging, primarily due to large acoustic variability and limited availability of training data. While supervised fine-tuning of adult pre-trained models has shown promise, it often fails to capture group-specific characteristics variations among children. To address this, we introduce GRoup-Aware PARtial model Merging (GRAPAM), a parameter-efficient approach that combines unsupervised clustering, partial fine-tuning, and model merging. Our approach adapts adult-pre-trained models to children by first grouping the children’s data based on acoustic similarity. Each group is used to partially fine-tune an adult pre-trained model, and the resulting models are merged at the parameter level. Experiments conducted on the MyST children’s speech corpus indicate that GRAPAM achieves a relative improvement of 6% of Word Error Rate (WER) , using the same amount of data, outperforming full fine-tuning while training fewer parameters. These results highlight the promise of model merging as a scalable and effective strategy for children’s ASR.
I Introduction
Recent progress in Automatic Speech Recognition (ASR) has been achieved through the introduction of large-scale Transformer-based architectures. These models have achieved state-of-the-art performance on numerous adult speech corpora, driven primarily by two key factors: the substantial increase in model complexity, reaching billions of parameters in some configurations [radford2023robust, hsu2021hubert, peng25c_interspeech], and the availability of large training datasets, with some exceeding one million hours of speech [koluguri2025granary]. Despite these advancements, using state-of-the-art ASR systems for children’s speech remains challenging, with performance consistently lagging behind that achieved for adult speech [sri_end2end, gelin2021endtoend, rolland2024exploring]. This disparity can be mainly attributed to the high acoustic variability characteristic of children’s speech, both between and within individual speakers. These variabilities primarily arise from the development of the child’s speech production system, leading to fluctuations in fundamental frequency, shifting of formant structures, and substantial variation in both spectral and temporal features of speech [Acoustic_change_children, reviewASRchildren]. In addition, children’s limited phonetic and linguistic maturity further impedes accurate speech recognition [language_children2], posing an additional challenge for existing ASR systems.
An additional challenge is the limited availability of large-scale and diverse children’s speech datasets, which significantly hinders the development of robust ASR models trained from scratch [sri_end2end, gelin2021endtoend]. In response, prior works have investigated various data augmentation strategies designed to expand and diversify existing corpora. These include techniques such as pitch normalisation [pitchnorm, yadav2019significance], Vocal Tract Length Normalisation (VTLN)[VTLN2, patel2024improving], multi-task learning [rolland2022multilingual], and the incorporation of synthetically generated speech data [wang2021towards, kadyan2021synthesis, rolland2024improved, zhao23c_interspeech].
A widely adopted approach for improving children’s ASR performances is Supervised Fine-Tuning (SFT) [sri_end2end, gelin2021endtoend, fan2022towards, fan2024benchmarking, jain2023adaptation]. In this approach, the parameters of a pre-trained ASR model, originally trained on a large adult speech corpus are adapted to children’s speech via additional supervised training. By leveraging the general linguistic and acoustic representations learned during pretraining, SFT facilitates the effective transfer of knowledge to the children’s speech domain. Importantly, prior studies have demonstrated that it is often unnecessary to adapt the entire model, as fine-tuning only a subset is often sufficient to achieve comparable performance improvements [rolland2024introduction]. One of the principal advantages of SFT lies in its ability to achieve substantial gains in recognition accuracy with relatively limited training data, making it particularly suitable for low-resource child speech settings.
@@ I Introduction
Nonetheless, a significant challenge of SFT for children’s ASR lies in the substantial acoustic variation across different age groups [1255448, gerosa2009review]. As children’s speech undergoes developmental changes with age, applying a uniform fine-tuning strategy across all age ranges may fail to capture age-specific acoustic characteristics effectively. Recent findings support the use of age-specific ASR models, which consistently yield higher recognition accuracy [lu2022improving, lilles2025fine, hamalainen2014improving]. Howevever, as there is one model per group, these models typically require prior knowledge of the speaker’s age, which is not always available, and raise concerns surrounding the scalability and parameter efficiency of storing and managing multiple distinct model copies.
@@
Recently, model merging has emerged as a lightweight way to combine specialised models trained on different domains without joint retraining, particularly in the domain of large language models (LLMs) [yang2024model, wortsman2022model, yu2024language, jin2022dataless]. A key insight from recent work on model merging for LLMs is that the resulting merged model often preserves the capabilities of the individual constituent models [yu2024language, jin2022dataless]. While parameter-level model merging remains a relatively novel area of research, particularly within LLMs and Vision Language Models (VLMs) [chen2025bring, biggs2024diffusion], emerging studies have begun to explore its potential applicability in ASR [ducorroy25_interspeech, tan2021novo], including promising developments on children’s speech [shankar2025selective].
@@
    We demonstrate consistent improvements over conventional fine-tuning for children’s ASR.
@@
II Related work
II-A Model merging
In recent years, LLMs and VLMs have demonstrated improved capabilities in tasks such as zero-shot inference and general-purpose understanding [radford2019language, brown2020language]. Owing to their training on diverse data, these models exhibit significant flexibility and generalisation capabilities. Nonetheless, several challenges persist. These include the generation of harmful or biased content [navigli2023biases], substantial high computational demands [10756602] and financial costs associated with training [li2023flm], and underperformance on specialised tasks [anjum2025domain].
@@ II-A Model merging
In this context, model merging has emerged as a promising research direction to overcome these limitations. It aims to combine multiple task-specific models into a single unified architecture that preserve the capabilities of each constituent model [wortsman2022model, matena2022merging, ilharco2022editing]. Unlike traditional multi-task learning, which relies on joint optimisation over multiple datasets, model merging operates directly at the parameter level, enabling the integration of pre-trained models without access to the original training data or additional retraining.
A variety of model merging strategies have been introduced in recent literature. Average Merging, or Linear Interpolation (LERP) [wortsman2022model], constructs a merged model by directly averaging the parameters of multiple fine-tuned models. Task Arithmetic [ilharco2022editing] computes task-specific vectors by subtracting a shared base model from each task-specific model; these vectors are then combined using predefined scaling factors to control the relative contribution of each task. Fisher Merging [matena2022merging] performs weighted parameters fusion, where the weights are derived from the Fisher Information Matrix, emphasising parameters that are more informative. RegMean [jin2022dataless] formulates the merging process as a linear regression problem with a closed-form solution. More recently, TIES-Merging [yadav2023ties] was introduced to mitigate parameter conflicts inherent in task arithmetic by trimming low-magnitude parameters, resolving sign inconsistencies, and separately merging parameters with consistent signs. Finally, DARE Merging [yu2024language] can be applied in conjunction with other methods, randomly dropping a proportion of parameters and rescaling the remaining weights prior to merging to improve robustness.
@@
Beyond the language and vision domains, model merging has also been explored in the context of ASR [ducorroy25_interspeech, tan2021novo]. An early approach, Divide-and-Merge (DAM) [tan2021novo], trained multiple models on different subsets of the dataset and subsequently combined them using genetic algorithms and SGD-based fine-tuning to merge the resulting acoustic models. More recently, model merging has been successfully applied to Whisper-based models fine-tuned for dysarthric speech recognition, demonstrating improved generalisation in low-resource and long-form scenarios [ducorroy25_interspeech]. Finally, Selective Attention Merging [shankar2025selective] introduced a layer-wise approach that merges attention-layer via task vectors between models trained on adult and child speech, yielding notable performance gains for children’s ASR.
II-B Partial Fine-tuning
As models continue to increase in scale, their parameter counts have grown dramatically. Consequently, full SFT of such large models is often impractical, especially in low-resource settings. In many cases, fine-tuning the entire model leads to suboptimal performance, as the vast number of trainable parameters can easily result in overfitting when only limited adaptation data is available. To address this challenge, parameter-efficient fine-tuning (PEFT) methods have emerged as compelling alternatives. Techniques such as Adapters [houlsby2019parameter, fan22d_interspeech, 10447091] and Low-Rank Adaptation (LoRA) [hu2021loralowrankadaptationlarge, 10447004] introduce additional trainable parameters while keeping the original pre-trained model weights frozen, thus enabling efficient and scalable adaptation. Another related approach, known as Partial Fine-Tuning (PFT) [ye2023partialfinetuningsuccessorfinetuning, Shen_Liu_Qin_Savvides_Cheng_2021, rolland2024exploring], involves updating only a subset of the existing parameters, typically within selected layers or components such as attention mechanisms or feed-forward networks, while leaving the rest of the model unchanged.
III Group-Aware Partial Model Merging
@@
Figure 1: Overview of the four stages of the GRAPAM pipeline. With $\mathcal{D}$ the full dataset and $\theta_{0}$ the pre-trained adult model parameters. The fire icon denotes stages involving ASR fine-tuning.
@@
In this work, we introduce GRAPAM, a novel approach inspired by several complementary lines of research: the demonstrated benefits of age-dependent ASR models [lu2022improving, lilles2025fine, hamalainen2014improving], the emerging success of model merging in children’s ASR [shankar2025selective], the divide-and-merge paradigm introduced in DAM [tan2021novo] and recent advances in partial fine-tuning techniques [rolland2024introduction]. GRAPAM is designed to improve performance on children’s speech by addressing speaker variability in a parameter-efficient manner through model merging. The method comprises four core stages, as illustrated in Figure 1.
Let $\mathcal{D}$ denote the training dataset of size $N$, defined as $\mathcal{D}={(x_{i},y_{i})}_{i=1}^{N}$, where $x_{i}$ represents the input speech signal and $y_{i}$ denotes the corresponding transcription.
@@
First, $\mathcal{D}$ is partitioned into similarity-based groups. Although age would be a natural criterion, in most children’s datasets, age information is not available. Moreover, chronological age may not accurately reflect developmental maturity. To address this, we adopt an unsupervised clustering strategy inspired by prior work [rolland2024exploring]. We extract utterance embeddings $E=\{e_{1},e_{2},\ldots,e_{N}\}$, where a single vector represent each utterance. In a second stage, we apply a clustering algorithm to divide these embeddings into K groups. A standard clustering algorithm such as k-means may be employed for this purpose:
$\min_{\{C_{1},\ldots,C_{K}\}}\sum_{i=1}^{K}\sum_{x\in C_{i}}\|x-\mu_{i}\|^{2},$ (1)
@@
with $\mathcal{M}\subseteq\{\theta_{all},\theta_{1},\dots,\theta_{k}\}$ the set of selected fine-tuned models for merging. While we use a uniform weight merge strategy, we note that $\alpha_{i}$ could be further optimised based on validation performance in future work. Finally, the resulting merged model, with parameters denoted as $\theta_{\text{GRAPAM}}$, is used for inference on the entire children’s test set. This unified model is expected to more effectively capture speaker-group-specific characteristics, thereby improving recognition performance.
IV Experimental setup
IV-A Corpus
TABLE I: My Science Tutor Children Speech Corpus statistics
@@ IV-A Corpus
For our experiments, we used the My Science Tutor Corpus (MyST) dataset. MyST is one of the largest publicly accessible collections of English children’s speech, comprising approximately 400 hours. It encompasses dialogues between children and a virtual tutor across eight scientific domains, involving 1,372 students in grades three to five. The corpus is pre-partitioned, ensuring equitable representation of scientific domains and unique student occurrences within each partition. However, only 45% of utterances are transcribed at the word level. In this work, we excluded utterances shorter than one second, mainly containing silence, and those longer than 20 seconds, due to GPU constraints. After this filtering process, the dataset comprises 56,859 utterances from 738 speakers, totaling approximately 154 hours of speech. More detailed statistics of the different partitions are provided in Table I.
IV-B Implementation details
@@
All experiments were conducted using the SpeechBrain toolkit [ravanelli2021speechbrain]. For the ASR model, we employed Whisper-medium [^1] https://huggingface.co/openai/whisper-medium.en [radford2023robust], which has previously demonstrated strong performance when fine-tuned on children’s speech [jain23_interspeech]. Whisper-medium is a Transformer-based encoder-decoder architecture comprising approximately 763.9M parameters, with 24 encoder layers and 24 decoder layers. The model was originally trained on a large-scale, multilingual dataset consisting of approximately 680k hours of transcribed speech.
In this work, we evaluate three types of utterance-level representations for clustering. First, we extract 192-dimensional speaker embeddings using a pre-trained ECAPA-TDNN model [^2] https://huggingface.co/speechbrain/spkrec-ecapa-voxceleb. Second, we compute a set of 100-dimensional low-level speech metrics using the Librosa library [mcfee2015librosa], including pitch, pause rate, speaking rate, spectral centroid, and signal-to-noise ratio; these features are subsequently reduced to 16 dimensions using Principal Component Analysis (PCA). Third, we introduce a one dimensional WER-based representation derived from zero-shot inference WERs using the Whisper model, clustering utterances based on their predicted transcription quality.
For all three representation types, we apply the $k$-means clustering algorithm from scikit-learn [pedregosa2011scikit], with a fixed number of clusters. We observed that, despite employing k-means, the cluster sizes remain balanced across all methods. Unless stated otherwise, we use $K=3$ clusters throughout our experiments. As for comparison, we also include random clustering, which assigns utterances to clusters uniformly at random.
For Partial Fine-Tuning (PFT), we experiment with three configurations: fine-tuning (i) all model parameters (763.9M), (ii) only the attention components (302.4M), and (iii) only the feed-forward network (402.9M) (FFN) components across all Transformer layers. All experiments are conducted on a single NVIDIA RTX A6000 GPU (48 GB), with a batch size of 16, a learning rate of $1\times 10^{-5}$, and one epoch of training. The models are optimised using the negative log-likelihood (NLL) loss function.
Finally, for model merging, we used LERP with uniform interpolation of the different models present in the configuration, ensuring that each group of children contributed equally to the final merged model.
V Results and discussion
To comprehensively assess the effectiveness of GRAPAM, we conduct a series of ablation studies aimed at disentangling the contributions of its core components. In particular, we evaluate the impact of different utterance clustering strategies, the effect of varying the number of groups, the number of selected merged models, the role of PFT, and alternative merging strategies including iterative model merging. These experiments are presented and analysed in detail in the remainder of this section.
It is important to note that this study does not explore the influence of model size, using self-supervised learning (SSL) models instead of Whisper, or the volume of training data. These factors have already been examined in prior work on model merging for children’s ASR [shankar2025selective].
V-A Group-aware model merging
@@
TABLE II: Results of different combinations for GRAPAM evaluated in WER (%). Values highlighted in grey indicate improvement relative to full-model fine-tuning on the entire dataset. The best-performing combinations are shown in bold.
$\mathcal{M}$ Clustering method
$\theta_{all}$ $\theta_{1}$ $\theta_{2}$ $\theta_{3}$ Spk-emb Random ZS WER LSM
14.05
x 9.95
x 10.34 10.11 10.25 9.73
x 9.79 9.69 10.32 10.61
x 10.36 10.63 10.18 10.03
x x 9.48 9.91 10.52 10.12
x x 9.77 10.34 10.09 9.45
x x 9.41 9.74 9.88 9.84
x x x 9.68 9.68 10.01 9.41
x x 9.72 9.98 10.32 9.79
x x 9.41 9.67 9.86 9.83
x x 9.76 9.77 9.96 9.70
x x x 9.36 9.64 9.98 9.97
x x x 9.68 9.98 9.72 9.68
x x x x 9.65 9.65 9.93 9.63
Average 9.59 9.84 10.03 9.74

### v2

# Rolland Abad

@@ Abstract
While supervised fine-tuning of adult pre-trained models for children's ASR has shown promise, it often fails to capture group-specific characteristics and variations among children. To address this, we introduce GRoup-Aware PARtial model Merging, a parameter-efficient approach that combines unsupervised clustering, partial fine-tuning, and model merging. Our approach adapts adult-pre-trained models to children by first grouping the children's data based on acoustic similarity. Each group is used to partially fine-tune an adult pre-trained model, and the resulting models are merged at the parameter level. Experiments conducted on the MyST children's speech corpus indicate that GRAPAM achieves a relative WER improvement of 6%, using the same amount of data, outperforming full fine-tuning while training fewer parameters.
keywords:
Children's speech, Children's ASR, Model merging, ASR, Partial fine-tuning
1 Introduction
Recent advances in Automatic Speech Recognition (ASR) have been made possible by the use of large-scale models, with state-of-the-art results on adult speech corpora [radford2023robust, hsu2021hubert, peng25c_interspeech]. However, performance on children’s speech remains consistently inferior to that on adult speech [sri_end2end, gelin2021endtoend, rolland2024exploring], primarily due to the pronounced acoustic variability of children’s productions both inter and intra-speaker, arising from ongoing physiological and articulatory development that affects fundamental frequency, formant patterns, and broader spectral–temporal characteristics [Acoustic_change_children, reviewASRchildren]. This difficulty is exacerbated by children’s limited phonetic and linguistic maturity [language_children2]. In addition, progress is further constrained by the scarcity of large, diverse children’s speech corpora, which limits the feasibility of training robust models from scratch [sri_end2end, gelin2021endtoend]. Consequently, existing work has explored augmentation and adaptation strategies such as pitch normalisation [pitchnorm, yadav2019significance], Vocal Tract Length Normalisation (VTLN) [VTLN2, patel2024improving], multi-task learning [rolland2022multilingual], and the use of synthetic speech [wang2021towards, kadyan2021synthesis, rolland2024improved, zhao23c_interspeech]. Among these, supervised fine-tuning (SFT) arise as a widely adopted strategy [sri_end2end, gelin2021endtoend, fan2022towards, fan2024benchmarking, jain2023adaptation], adapting an adult-pretrained ASR model to children’s speech via additional training and thereby transferring general acoustic–linguistic representations to children's specifics. Importantly, it was observed that comparable gains can often be obtained by updating only a subset of parameters rather than the full network [rolland2024introduction], making SFT particularly effective in low-resource child-speech settings where substantial accuracy improvements are achievable with limited data.
@@ 1 Introduction
Nonetheless, a significant challenge of SFT for children's ASR lies in the substantial acoustic variation across different age groups [1255448, gerosa2009review]. As children’s speech undergoes developmental changes with age, applying a uniform fine-tuning strategy across all age ranges may fail to capture age-specific acoustic characteristics effectively. Recent findings support the use of age-specific ASR models, which consistently yield higher recognition accuracy [lu2022improving, lilles2025fine, hamalainen2014improving]. However, group-specific fine-tuning produce many checkpoints raising concerns surrounding the scalability and parameter efficiency of storing and managing multiple distinct model copies. Additionally, these models typically require prior knowledge of the speaker’s age, which is not always available, making this group-specific training unavailable therefore losing its associated benefits.
@@
Model merging has emerged as a alternative way to combine specialised models trained on different domains without joint retraining, particularly in the domain of large language models (LLMs) [yang2024model, wortsman2022model, yu2024language, jin2022dataless]. A key insight from recent work on model merging for LLMs is that the resulting merged model often preserves the capabilities of the individual constituent models [yu2024language, jin2022dataless]. While parameter-level model merging remains a relatively novel area of research, particularly within LLMs and Vision Language Models (VLMs) [chen2025bring, biggs2024diffusion], emerging studies have begun to explore its potential applicability in ASR [ducorroy25_interspeech, tan2021novo], including promising developments on children’s speech [shankar2025selective].
@@
    We demonstrate consistent improvements over conventional fine-tuning for children's ASR.
@@
2 Related work
2.1 Model merging
@@
Model merging has emerged as a promising research direction to combine multiple task-specific models into a single unified architecture that preserve the capabilities of each constituent model [wortsman2022model, matena2022merging, ilharco2022editing]. Unlike traditional multi-task learning, which relies on joint optimisation over multiple datasets, model merging operates directly at the parameter level, enabling the integration of pre-trained models without access to the original training data or additional retraining.
Recent work has proposed diverse model merging strategies. Average Merging, or Linear Interpolation (LERP) averages parameters across fine-tuned models [wortsman2022model], while Task Arithmetic combines task vectors formed by subtracting a shared base model [ilharco2022editing]. Fisher Merging uses Fisher-information weights to emphasise informative parameters [matena2022merging], and RegMean casts merging as closed-form linear regression [jin2022dataless]. More recent methods address interference and robustness: TIES trims small-magnitude updates and resolves sign conflicts before merging [yadav2023ties], and DARE randomly drops and rescales parameters as a pre-merge regulariser [yu2024language].
@@
Model merging has also been explored for ASR [ducorroy25_interspeech, tan2021novo]. Divide-and-merge (DAM) trains models on data partitions and combines them via genetic search and SGD fine-tuning [tan2021novo]. Recent studies apply merging to Whisper checkpoints for dysarthric speech, improving generalisation in low-resource and long-form settings [ducorroy25_interspeech], and propose Selective Attention Merging, a layer-wise approach that merges attention layers via task vectors for adult–child transfer, improving children’s ASR [shankar2025selective].
2.2 Partial Fine-tuning
As model scale increases, full SFT becomes increasingly costly and, in low-resource regimes, can degrade performance due to overfitting from updating hundreds of millions of parameters. Parameter-efficient fine-tuning (PEFT) addresses this by adapting models with limited trainable parameters: adapter-based methods [houlsby2019parameter, fan22d_interspeech, 10447091] and LoRA [hu2021loralowrankadaptationlarge, 10447004] add small trainable modules while keeping the backbone frozen. A closely related strategy, partial fine-tuning (PFT) [ye2023partialfinetuningsuccessorfinetuning, Shen_Liu_Qin_Savvides_Cheng_2021, rolland2024exploring], instead updates only selected subsets of the original parameters (e.g., attention or feed-forward blocks), leaving the remainder unchanged.
3 Group-Aware Partial Model Merging
@@
Figure 1: Overview of the four stages of the GRAPAM pipeline. With $\mathcal{D}$ the full dataset and $\theta_{0}$ the pre-trained adult model parameters. The fire icon denotes stages with ASR fine-tuning.
@@
We introduce GRAPAM, a novel method that combines insights from several complementary research directions: prior evidence for the effectiveness of age-dependent ASR models [lu2022improving, lilles2025fine, hamalainen2014improving], recent progress on model merging for children’s ASR [shankar2025selective], the divide-and-merge paradigm of DAM [tan2021novo], and advances in partial fine-tuning for parameter-efficient adaptation [rolland2024introduction]. GRAPAM targets children’s speech by mitigating speaker related variability using parameter-efficient fine-tuning followed by model merging. The approach consists of four stages, illustrated in Figure 1. Let $\mathcal{D}$ denote the training dataset of size $N$, defined as $\mathcal{D}={(x_{i},y_{i})}_{i=1}^{N}$, where $x_{i}$ represents the input speech signal and $y_{i}$ denotes the corresponding transcription.
@@
First, $\mathcal{D}$ is partitioned into similarity-based groups. Although age would be a natural criterion, in most children's datasets, age information is not available. Moreover, chronological age may not accurately reflect developmental maturity. To address this, we adopt an unsupervised clustering strategy inspired by prior work [rolland2024exploring]. We extract utterance embeddings $E=\{e_{1},e_{2},\ldots,e_{N}\}$, where a single vector represent each utterance. In a second stage, we apply a clustering algorithm to divide these embeddings into K groups. A standard clustering algorithm such as k-means may be employed for this purpose:
$\min_{\{C_{1},\ldots,C_{K}\}}\sum_{i=1}^{K}\sum_{x\in C_{i}}\lVert x-\mu_{i}\lVert^{2}$ (1)
@@
with $\mathcal{M}\subseteq\{\theta_{all},\theta_{1},\dots,\theta_{k}\}$ the set of selected fine-tuned models for merging. While we use a uniform weight merge strategy, we note that $\alpha_{i}$ could be further optimised based on validation performance in future work. Finally, the resulting merged model, with parameters denoted as $\theta_{\text{GRAPAM}}$, is used for inference on the entire children’s test set.
4 Experimental setup
4.1 Corpus
Table 1: My Science Tutor Children Speech Corpus statistics
@@ 4.1 Corpus
For our experiments, we use the My Science Tutor Corpus (MyST) dataset, one of the largest publicly available collections of English children’s speech ($\sim$400 hours). MyST contains child–virtual tutor dialogues spanning eight science domains and includes recordings from 1,372 students in grades 3–5. The dataset is released with predefined partitions that balance scientific domains and ensure that each student appears in only one split. Only 45% of utterances have word-level transcripts.
We discard utterances shorter than 1 s, dominated by silence, and those longer than 20 s due to GPU constraints. After filtering, the resulting dataset contains 56,859 utterances from 738 speakers, totaling $\sim$154 hours of speech. Detailed partition statistics are reported in Table 1.
4.2 Implementation details
@@
All experiments were performed with SpeechBrain [ravanelli2021speechbrain] using Whisper-medium.en [radford2023robust], a 24-layer encoder–decoder Transformer with 763.9M parameters, pre-trained on $\sim$680k hours of multilingual speech and shown effective for children’s ASR [jain23_interspeech].
We cluster utterances using three representation types: (i) 192-d speaker embeddings from a pre-trained ECAPA-TDNN model [^1] https://huggingface.co/speechbrain/spkrec-ecapa-voxceleb, (ii) 16-d acoustic descriptors obtained by extracting 100 Librosa low-level metrics [mcfee2015librosa] and applying PCA, and (iii) a 1-d WER-based score from Whisper zero-shot predictions. For each representation, we apply k-means from scikit-learn [pedregosa2011scikit] (default K=3); we also report a random clustering baseline.
For PFT, we compare updating (i) all parameters, (ii) attention modules only, or (iii) FFN modules only. Training uses a single RTX A6000 (48GB), batch size 16, learning rate $10^{-5}$, one epoch, and NLL loss. For merging, we use LERP with uniform weights so each cluster contributes equally to the merged model.
5 Results
@@
Table 2: WER (%) results for GRAPAM. Left: clustering-method variants. Right: partial fine-tuning variants (SFT/PFT). Grey indicates improvement relative to full-model fine-tuning on the entire dataset; diagonal pattern indicates improvement over full SFT but not over the corresponding PFT on $\mathcal{D}$. Best values are in bold.
$\mathcal{M}$ Clustering method Fine-tuning method
$\theta_{all}$ $\theta_{1}$ $\theta_{2}$ $\theta_{3}$ Spk-emb Random ZS WER LSM Full SFT PFT FFN PFT ATTN
14.05 014.05
x 09.95 9.95 9.48 9.46
x 10.34 10.11 10.25 9.73 10.34 10.03 10.00
x 9.79 9.69 10.32 10.61 9.79 9.68 9.68
x 10.36 10.63 10.18 10.03 10.36 9.58 9.71
x x 9.48 9.91 10.52 10.12 9.48 9.52 9.61
x x 9.77 10.34 10.09 9.45 9.77 9.69 9.68
x x 9.41 9.74 9.88 9.84 9.41 9.44 9.56
x x x 9.68 9.68 10.01 9.41 9.68 9.45 9.60
x x 9.72 9.98 10.32 9.79 9.72 9.66 9.66
x x 9.41 9.67 9.86 9.83 9.41 9.32 9.62
x x 9.76 9.77 9.96 9.70 9.76 9.66 9.66
x x x 9.36 9.64 9.98 9.97 9.36 9.33 9.68
x x x 9.68 9.98 9.72 9.68 9.68 9.32 9.66
PFT x x x - - - - - 9.36 9.51
x x x x 9.65 9.65 9.93 9.63 9.65 9.31 9.38
Average 9.59 9.84 10.03 9.74 9.59 9.47 9.60
5.1 Group-aware model merging